Step 1 of 2 : Suppose a sample of 370 tankers is drawn. Of these ships, 311 did not have spills. Using the data, estimate the proportion of oil tankers that had spills. Enter your answer as a fraction or a decimal number rounded to three decimal places.
Answer #1

Solution: Given that n = 370, x = 311

Step 1 of 2 : q = x/n = 311/370 = 0.841

The proportion of oil tankers that had spills : 1 - q = 1-0.841 = 0.159